A certain sum is divided among A, B and C in a manner that for every rupee A gets, B gets 60 paise, and that for every rupee B gets, C gets 55 paise. If C's share in the total sum is Rs 66, then what is the share of A?
Answer by mananth(16946)


A certain sum is divided among A, B and C in a manner that for
every rupee A gets, B gets 60 paise, so the ratio is 1:0.6
a/b = x/1
x=1/0.6 = 1.66
and that for every rupee B gets, C gets 55 paise. the ratio 1:0.55
If C's share in the total sum is Rs 66,
the ratio will be 1.66 : 1: 0.55
If C gets 66 B will get 66/0.55 =120
A will get 120*1.66 =199.2
199.2:120:66